11. Given the points below, find XY. Round to

the nearest hundredth.

X1-9, 2) and YI5, -4)


1
Expert's answer
2021-10-27T13:40:06-0400

"|XY|=\sqrt{(5+9)^2+(-2-4)^2}=\sqrt{232}\approx 15.23."